# BEADs: Bias Evaluation Across Domains

### v3

@@ 5.4 Language Generation Task
Task Setting: This task generates neutral or benign versions of biased texts, the goal is to remove bias in language while retaining knowledge or context of the original sentence. The dataset portion has 8,300 entries with various demographics and bias aspects. Format: Each sentence pair is presented as <original sentence>, <benign variation>. For example, Original Sentence: “She lacks the aggressive approach needed for effective leadership in this role", Benign Sentence: “The candidate’s leadership style does not align with the assertive approach required for this role". The dataset of our dataset for this purpose is Language Generation Dataset. We also format this dataset portion into Alpaca [28] format for instruction-fine-tuning.
@@ 6 Discussion
In this work, we used a two-phase approach: initially, we employed GPT-3.5 for preliminary labeling and then utilized GPT-4 for creating higher-quality, gold-standard labels. This choice was driven by the experimental nature of our work and the availability of newer model versions, which encouraged us to explore the latest advancements. However, we also acknowledge that the associated costs of using these models can be quite high and may not always be feasible for everyone.
@@ 7 Conclusions
In this work, we present BEADs, a comprehensive dataset specifically labeled for bias detection across multiple NLP tasks, including text classification, token classification, bias quantification, and debiasing through benign language generation. The dataset is annotated using a two-phase approach that combines active learning with GPT-3.5-generated labels and higher-quality labels provided by GPT-4, which are then verified by expert annotators for reliability. Benchmarking the dataset with BERT-like models and autoregressive LLMs, we find that smaller models, such as BERT and RoBERTa, outperform LLMs in bias classification tasks. Furthermore, we observe that some LLMs tend to perpetuate biases for certain demographics, with smaller language models being more prone to biases. However, fine-tuning LLMs on safe data from BEADs tends to reduce these biases while maintaining the models’ overall performance.
Acknowledgements
We extend our gratitude to the Province of Ontario, the Government of Canada through CIFAR, and the corporate sponsors of the Vector Institute for their generous support and provision of resources essential for this research. Further details on our sponsors can be found at www.vectorinstitute.ai/#partners. We also acknowledge our expert review team and everyone involved in the data review process.
CRediT authorship contribution statement
Shaina Raza: Conceptualization, Investigation, Formal analysis, Methodology, Project administration, Software, Experimentation, Validation, Visualization, Writing – original draft, Writing – review, & editing, Supervision. Mizanur Rahman: Investigation. Experimentation. Formal Analysis. Validation. Review. Michael R. Zhang : Validation, Writing – review& editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data and code is made available.
Appendix
$\bigtriangleup$ Warning: This paper contains examples that may be considered offensive.
Appendix A Comparison with Existing Datasets
Table 9: Comparison of Bias Datasets: Each entry includes the dataset name (Dataset), source (Source), dataset details, annotation style (Annot.: Human (H) or Model-based (M)), accessibility (Size/Annot./Open), primary task (Task), and bias type (Dimension). Acronyms: O: Open access (✓for ‘yes’), H: human-labeled, quant.: quantification, eval.: evaluation, LM: language model, H-M: Human & Model-based annotations.
Dataset Source Size/Annot./ Open Task Dimension Data Format Year
Racist [77] Twitter 16.9K/H/✓ Classification Hate speech NLP 2016
HSOL [78] Twitter 24.8K/H/✓ Classification Hate speech NLP 2017
Winogender [36] Winograd schemas 720/H/✓ Coreference resolution Gender bias NLP 2018
WinoBIAS [50] Winograd schemas 3.2K/H/✓ Coreference resolution Gender bias NLP 2018
White Supremacy [79] StormFront 9.9K/H/✓ Classification Hate speech NLP 2018
Toxic Comment [12] Wikipedia 159.6K/H/✓ Classification Toxicity NLP 2018
Age Bias [22] Blogs, Tweets 4.2K blogs; 64.3K comments; 13.8K tweets/H-M/✓ Classification Ageism, sentiment NLP 2018
Twitter Abusive [80] Twitter 80K/H/na Classification Hate NLP 2018
Jigsaw [13] Twitter 1.78M/H/✓ Classification Toxicity NLP 2019
Hyperpartisan [62] News publishers 1.3K/H; 754K/✓ Classification Politics NLP 2019
OLID [81] Twitter 14K/H/✓ Classification Toxicity NLP 2019
BASIL [82] FOX/NYT/ HPO 300/H/✓ Classification Political bias NLP 2019
Microaggressions [83] Reddit 2.9K/H/✓ Classification Hate speech NLP 2019
Wiki Neutrality [84] Wikipedia 565K edits/✓ Token classification Framing, epistemological biases NLP 2020
Ukraine Crisis [63] URLs 2.1K/H/✓ Analysis Political bias NLP 2020
SOLID [85] Twitter 9M/Model/ Request Classification Offensive NLP 2020
Realtoxicity Prompts [17] Webtext corpus 100K/Model/✓ Classification Toxicity Prompts 2020
Stereotype [16] Wikidata 17K/H/✓ Bias quant. Gender, profession, race NLP 2020
SBIC [64] Reddit, Twitter 44.7K/H/✓ Classification Hate speech NLP 2020
DynaHate [86] Mixed 41.3K/H/✓ Classification Hate speech NLP 2020
RedditBIAS [18] Reddit 10.9K/H/✓ LM eval. Religion, race, gender, queerness NLP 2021
HateXplain [51] Twitter 20.1K/H/✓ Classification Hate speech NLP 2021
Sexist [87] Twitter 16.3K/H/✓ Classification Sexism NLP 2021
UsVsThem [88] Reddit 6.9K/H/Request Classification Political bias NLP 2021
HateCheck [89] Research, Interviews 3.7K/H/✓ Classification Hate NLP 2021
BOLD [15] Wikipedia 23.7K prompts/H/✓ Classification Profession, gender, race Prompt 2021
BBQ [37] Human-written 58K/H/✓ QA Age, disability status, gender identity, nationality, physical appearance, race/ethnicity, religion, socioeconomic status, sexual orientation Prompts 2021
HateMoji [90] Mixed Hatemoji-check 3.9K; Hatemoji-build 5.9K/H/✓ Classification Hate speech NLP 2022
BIGNEWS [91] Media outlets 3.69M/ Algorithmic/✓ Prediction Political bias NLP 2022
Toxigen [21] GPT-3 274K/Human (9.9k) and machine (274K)/✓ Classification Implicit hate speech NLP, prompts 2022
HolisticBias [14] Hand-crafted 460K prompts/H/✓ Bias quant. Sociodemographics, bias: ability, age, body type, characteristics, cultural, gender/sex, nationality, nonce, political, race/ethnicity, religion, sexual orientation, socioeconomic Prompts 2022
MBIB [73] Various datasets 4.24M/pre-labeled/✓ Classification Multiple biases NLP 2023
SeeGULL [92] Hybrid 7.75K/H-M/✓ Bias quant. Geo-political Categorical 2023
HolisticBiasR [19] RealToxicity Prompts, BOLD, Toxigen v2, AdvPromptSet, Regard dataset 214.46K/previous/✓ Bias quant. Age, body type, class, culture, disability, gender, nationality, political ideology, race/ethnicity, religion, sexual orientation Prompts 2023
WinoQueer [38] Hybrid 45K/H/✓ Bias quant. Gender, sexual orientation NLP 2023
HarmfulQ [93] Prompts 1.9K/Prompt-based/✓ QA/LLM safety Harmful conversations Prompt 2023
BEADs Multiple benchmarks and scraping 3.67M/H-M/✓ Sentence classification (bias, toxicity, sentiment), token classification, Demographic groups, Bias quantification, language generation Bias, hate speech, toxicity, stereotypes NLP, instruction fine-tuning datasets, prompts -
Appendix B Prompts for Annotation
B.1 Prompt for Initial Data Labeling
In initial experiments, we manually inspected labels and found zero-shot and few-shot performance to be similar. For lower cost, we used zero-shot to do the labeling on the gold set.
You are a language model tasked with identifying the following elements in a given text:
1. Bias classification: Indicate if the text is highly biased, slightly biased, or neutral.
2. Toxicity: Indicate if the text is toxic or non-toxic.
3. Sentiment: Indicate if the sentiment of the text is positive, negative, or neutral.
4. Biased words: Terms that convey prejudice or stereotypes.
5. Aspects of bias: The context or topic where the bias appears.
6. Demographic mentions: References to specific groups based on attributes like gender, race, religion, age, disability, and sexual orientation.
Please analyze the following text and provide your response in the specified format.
Example 1: Text: “Women are not good at technical jobs and should stick to domestic tasks".
Bias classification: Highly biased
Toxicity: Non-toxic
Sentiment: Negative
Biased words: “not good at technical jobs", “should stick to domestic tasks“
Aspects of bias: Gender bias in professional capabilities
Demographic mentions: Women
Example 2: Text: “People with mental health issues are unreliable and cannot handle stress".
Bias classification: Highly biased
Toxicity: Non-toxic
Sentiment: Negative
Biased words: “unreliable", “cannot handle stress”
Aspects of bias: Bias against mental health conditions
Demographic mentions: People with mental health issues
Example 3 (Neutral Example): Text: “The conference was attended by professionals from various fields to discuss the latest advancements in technology".
Bias classification: Neutral
Toxicity: Non-toxic
Sentiment: Neutral
Biased words: None
Aspects of bias: None
Demographic mentions: Professionals from various fields
Your Turn: Text: Insert text here Bias classification: Toxicity: Sentiment: Biased words: Aspects of bias: Demographic mentions:
B.1.1 Labeling for Classification
@@
TEXT: “I think most startups fail because their founders are too inexperienced". BIAS LABEL: Biased TOXICITY LABEL: Non-toxic SENTIMENT LABEL: Negative NOTABLE WORDS/PHRASES: “too inexperienced“
TEXT: “Anyone who supports that policy is clearly an idiot". BIAS LABEL: Biased TOXICITY LABEL: Toxic SENTIMENT LABEL: Negative NOTABLE WORDS/PHRASES: “clearly an idiot“
TEXT: “It’s hard to say if there’s any truth to such claims, but we should investigate further". BIAS LABEL: Neutral TOXICITY LABEL: Non-toxic SENTIMENT LABEL: Neutral NOTABLE WORDS/PHRASES: N/A
TEXT: “All politicians are liars and only care about their own gain". BIAS LABEL: Biased TOXICITY LABEL: Toxic SENTIMENT LABEL: Negative NOTABLE WORDS/PHRASES: “All politicians are liars“
TEXT: “This charity does wonderful work for the community and has helped many people". BIAS LABEL: Neutral TOXICITY LABEL: Non-toxic SENTIMENT LABEL: Positive NOTABLE WORDS/PHRASES: “wonderful work", “helped many people“
B.1.2 Demographic Mention Annotation Task
@@
TEXT: “A young woman in her 20s, recently graduated from college, discussed the challenges faced by her generation in finding stable employment in today’s economy". Demographic Mentions:
  • 1.
    Age Mention: “young", “in her 20s", “her generation“
  • 2.
@@
TEXT: “The Muslim community in our city has contributed significantly to the local economy and culture". Demographic Mentions:
  • 1.
@@
  • 2.
    Ethnic Mention: Implicit mention through “community", suggesting a specific group’s cultural impact.
TEXT: “Veterans, especially those who served in recent conflicts, face unique challenges in re-adapting to civilian life". Demographic Mentions:
  • 1.
@@
TEXT: “Older adults are at a higher risk of severe illness from the flu, according to health experts". Demographic Mentions:
  • 1.
@@
B.1.3 Aspect of Bias Annotation Task
@@
TEXT: “Young people today have no real understanding of how to handle real-world problems". BIAS ASPECT: Age Bias NOTABLE WORDS/PHRASES: “Young people today", “no real understanding“
TEXT: “Men are typically more suited for high-stress jobs than women". BIAS ASPECT: Gender Bias NOTABLE WORDS/PHRASES: “Men", “more suited", “than women“
TEXT: “Despite their qualifications, immigrants from Asia often find it hard to get jobs in this sector". BIAS ASPECT: Ethnicity Bias NOTABLE WORDS/PHRASES: “immigrants from Asia", “hard to get jobs“
TEXT: “Atheists tend to have a more cynical view of life compared to those who are religious". BIAS ASPECT: Religious Bias NOTABLE WORDS/PHRASES: “Atheists", “more cynical view", “compared to those who are religious“
TEXT: “People from low-income neighborhoods are more likely to engage in criminal activities". BIAS ASPECT: Socioeconomic Bias NOTABLE WORDS/PHRASES: “low-income neighborhoods", “more likely to engage in criminal activities“
TEXT: “Western music is inherently more sophisticated than traditional music from Africa". BIAS ASPECT: Cultural Bias NOTABLE WORDS/PHRASES: “Western music", “more sophisticated", “than traditional music from Africa“
B.2 Initial Data Labeling and Active Learning
Step 1: Initial Label Generation with GPT-3.5
We used GPT-3.5 (gpt-3.5-turbo), equipped with few-shot prompts (2 examples), to generate initial labels for various features across over 50,000 samples. The labels generated are Biases, toxicity, sentiment, biased words, aspects, and demographic mentions.
Step 2: Refinement and Specific Task Labeling with BERT
The BERT model is fine-tuned on the initial labelled data for following tasks: - Text classification for bias, toxicity and sentiment.
- Token Classification - a multi-classification problem, trained for named entity recongition model based on BERT-large uncased for identifying specific entities within the text, such as bias words/ pharases.
- Identity Mentions: For identity mentions, we fine-tuned a BERT-large uncased model to identity these demographic mentions and then used this trained model to recognize and label demographic identities like gender and ethnicity from unlabelled data.
- Aspects Identification: We fine-tuned a BERT-large uncased model on aspects and used this trained model to find aspects of biases like gender, racism, sexism, lookism, abletism and such from the unlabelled data.
Step 3: Quality Control and Iterative Refinement
Outputs from BERT were reviewed, focusing on predictions with low confidence. Continuous improvement in data labeling accuracy and classifier performance through feedback and adjustment from our reviewers, details in Appendix C.
Step 4: Ensuring Consistency
Inter Annotator Agreement (IAA) is implemented in subsequent phases of the annotation process to validate the consistency and reliability of the labels across the dataset. We utilize Cohen’s Kappa statistic to determine IAA, which yields a substantial agreement level of 0.67.
In our work, we estimate the project duration to be approximately 185 weeks and the total expert effort to be about 5 hours per week, calculated based on processing 370 batches of data over this period with 20 reviewers.
Figure 4: Initial data labeling
B.3 Gold Data Labeling
The gold dataset creation began with a strategic sampling phase to ensure a balanced representation of classification labels and bias aspects. First, we filtered the 3.7M main dataset to obtain equal representation of labels from bias, toxicity, and sentiment classes, resulting in approximately 1.9M records. Next, we further filtered the dataset by aspects of bias, ensuring that each aspect appeared at least 50 times to maintain equal representation.
From this filtered dataset, we reduce the dataset to 50k records, ensuring equal representation from bias labels, toxicity labels, and sentiments. Each labeling is saved separately for training and testing. For the rest of the tasks, such as token classification, demographics and aspects, we used the bias labels - 50k sample.
Appendix C Dataset Reviewers
C.1 Guidelines
@@
C.2 Review Guidelines Checklist
@@
C.3 Annotation Performance Evaluation
The performance of GPT-4 as an annotator was evaluated based on three criteria: (1) accuracy, (2) consistency, and (3) coverage, using a subset of 500 manually labelled data. The table below summarizes the results:
Criterion Accuracy Consistency Coverage
Results 94% 85% 92%
Table 10: Performance Evaluation of GPT-4 as an Annotator
GPT-4 demonstrated high accuracy, correctly labeling 94% of the data points in the subset. Consistency was measured at 85%, indicating reliable performance across different types of data within the subset. The coverage metric, at 92%, shows GPT-4 capability to handle a wide range of topics and scenarios represented in the dataset.
Appendix D Initial Dataset Analysis and Results
The main dataset analysis is given in Table 11.
Table 11: Overview of the main dataset.
Attribute Details
Total Entries 3,674,927
Columns Text, Dimension, Biased Words, Aspect, Bias Label, Sentiment, Toxic, Identity Mention
Label Distribution Neutral: 1,886,253; Slightly Biased: 1,041,254; Highly Biased: 747,420
Sentiment Distribution Positive: 1,678,586; Negative: 1,224,531; Neutral: 771,810
Toxicity Classification Non-toxic: 2,448,078; Toxic: 1,226,849
Identity Mention Male: 152985, Female: 108042, Christian: 76366, White: 37262, Muslim: 23426, Black: 21701, Gay lesbian: 18795, Mental health condition: 14627, Jewish: 7749
Aspects 3496957 aspects with 1875 unique (Table LABEL:tab:aspects)
D.1 Text Classification
The main dataset classification labels are in Tale 12.
Table 12: Analysis of Label Distributions, Sentiments, and Toxicity
Label Type Label Count Sentiment Distribution (%) Toxicity (%)
Negative Neutral Positive Non-toxic Toxic
Bias Neutral 1,886,253 0.00 36.79 63.21 76.76 23.24
Slightly Biased 1,041,254 54.14 7.09 38.77 61.98 38.02
Highly Biased 747,420 88.40 0.53 11.06 47.48 52.52
Toxic Not Toxic 2,448,078 - - - 100.00 0.00
Toxic 1,226,849 - - - 0.00 100.00
Sentiment Positive 1,678,586 - - 100.00 - -
Negative 1,224,531 100.00 0.00 0.00 - -
Neutral 771,810 0.00 100.00 0.00 - -
The data and analysis in Table 12 shows that texts labeled as ‘Highly Biased’ are more likely to exhibit negative sentiments and higher toxicity levels, suggesting a correlation between strong bias and negative or inflammatory content. In contrast, texts classified as ‘Neutral’ generally show positive sentiments and lower toxicity, indicating less contentious content. ‘Slightly Biased’ articles display a more balanced sentiment distribution but still contain a significant proportion of toxic content, reflecting a nuanced portrayal of news and social media. These trends suggest that the degree of perceived bias in news and social media significantly influences both the sentiment and the toxicity of the content. Some examples are shown in Table 13.
Table 13: Examples of Texts with Bias, Toxicity, and Sentiment Labels
Text Bias Label Toxic Sentiment
Unbelievable incompetence leading to a recession! Highly Biased 1 Negative
Such ignorance in power should be criminal! Highly Biased 1 Negative
Blindly supporting that fool agenda is an intellectual failure. Highly Biased 1 Negative
The other party’s approach is always disastrous and corrupt. Highly Biased 1 Negative
Their proposal might work, but lacks depth ! go and do more work. Slightly Biased 1 Negative
Your point is valid, though I respectfully hold a different view. Slightly Biased 1 Neutral
An interesting perspective, albeit marred by some inaccuracies. Slightly Biased 1 Negative
Open to evidence-based discussion on this topic. Neutral 0 Neutral
My stance is neutral until more information emerges. Neutral 0 Negative
Innovations in clean energy are leading us to a brighter future. Neutral 0 Positive
Community efforts are making a visible difference in people’s lives. Neutral 0 Positive
Equal access to resources empowers societal growth and harmony. Neutral 0 Positive
Classification Results Evaluations on classification task of main data are in Table 14.
Table 14: Performance Metrics on the Main Data following 80-20 splits for Bias, Toxicity, and Sentiment
Model Accuracy Precision Recall
Bias
Bert-Large-Uncased 0.796 0.796 0.794
RoBERTa-Large 0.799 0.800 0.798
DistilBERT-based-Uncased 0.816 0.817 0.816
Electra-Base 0.759 0.761 0.759
GPT-2 Medium 0.697 0.697 0.696
GPT-2 Large 0.774 0.775 0.774
Toxicity
Bert-Large-Uncased 0.780 0.780 0.780
RoBERTa-Large 0.790 0.790 0.790
DistilBERT-based-Uncased 0.790 0.790 0.790
Electra-Base 0.760 0.760 0.760
GPT-2 Medium 0.815 0.815 0.815
GPT-2 Large 0.850 0.851 0.851
Sentiment
Bert-Large-Uncased 0.780 0.750 0.770
RoBERTa-Large 0.750 0.760 0.750
DistilBERT-based-Uncased 0.810 0.810 0.810
Electra-Base 0.710 0.730 0.710
GPT-2 Medium 0.820 0.829 0.824
GPT-2 Large 0.860 0.864 0.861
D.2 Aspect Categorization

### v4

@@ 5.4 Language Generation Task
Task Setting: This task generates neutral or benign versions of biased texts, the goal is to remove bias in language while retaining knowledge or context of the original sentence. The dataset portion has 8,300 entries with various demographics and bias aspects.
Format: Each sentence pair is presented as original sentence, benign variation. For example, Original Sentence: “She lacks the aggressive approach needed for effective leadership in this role”, Benign Sentence: “The candidate’s leadership style does not align with the assertive approach required for this role”. The dataset of our dataset for this purpose is Language Generation Dataset. We also format this dataset portion into Alpaca [24] format for instruction-fine-tuning.
@@ 6 Discussion
In this work, we used a GPT-4 for creating higher-quality, gold-standard labels. This choice was driven by the experimental nature of our work and the availability of newer model versions, which encouraged us to explore the latest advancements. However, we also acknowledge that the associated costs of using these models can be quite high and may not always be feasible for everyone.
@@ 7 Conclusions
In this work, we present BEADs, a comprehensive dataset specifically labeled for bias detection across multiple NLP tasks, including text classification, token classification, bias quantification, and debiasing through benign language generation. The dataset is annotated using GPT-4, which are then verified by expert annotators for reliability. Benchmarking the dataset with BERT-like models and autoregressive LLMs, we find that smaller models, such as BERT and RoBERTa, outperform LLMs in bias classification tasks. Furthermore, we observe that some LLMs tend to perpetuate biases for certain demographics, with smaller language models being more prone to biases. However, fine-tuning LLMs on safe data from BEADs tends to reduce these biases while maintaining the models’ overall performance.
Appendix A Prompts for Annotation
A.1 Labeling for Classification
@@
TEXT: “I think most startups fail because their founders are too inexperienced”. BIAS LABEL: Biased TOXICITY LABEL: Non-toxic SENTIMENT LABEL: Negative NOTABLE WORDS/PHRASES: “too inexperienced“
TEXT: “Anyone who supports that policy is clearly an idiot”. BIAS LABEL: Biased TOXICITY LABEL: Toxic SENTIMENT LABEL: Negative NOTABLE WORDS/PHRASES: “clearly an idiot“
TEXT: “It’s hard to say if there’s any truth to such claims, but we should investigate further”. BIAS LABEL: Neutral TOXICITY LABEL: Non-toxic SENTIMENT LABEL: Neutral NOTABLE WORDS/PHRASES: N/A
TEXT: “All politicians are liars and only care about their own gain”. BIAS LABEL: Biased TOXICITY LABEL: Toxic SENTIMENT LABEL: Negative NOTABLE WORDS/PHRASES: “All politicians are liars“
TEXT: “This charity does wonderful work for the community and has helped many people”. BIAS LABEL: Neutral TOXICITY LABEL: Non-toxic SENTIMENT LABEL: Positive NOTABLE WORDS/PHRASES: “wonderful work”, “helped many people“
A.2 Demographic Mention Annotation Task
@@
TEXT: “A young woman in her 20s, recently graduated from college, discussed the challenges faced by her generation in finding stable employment in today’s economy”. Demographic Mentions:
  • Age Mention: “young”, “in her 20s”, “her generation“
@@
TEXT: “The Muslim community in our city has contributed significantly to the local economy and culture”. Demographic Mentions:
@@
  • Ethnic Mention: Implicit mention through “community”, suggesting a specific group’s cultural impact.
TEXT: “Veterans, especially those who served in recent conflicts, face unique challenges in re-adapting to civilian life”. Demographic Mentions:
@@
TEXT: “Older adults are at a higher risk of severe illness from the flu, according to health experts”. Demographic Mentions:
@@
A.3 Aspect of Bias Annotation Task
@@
TEXT: “Young people today have no real understanding of how to handle real-world problems”. BIAS ASPECT: Age Bias NOTABLE WORDS/PHRASES: “Young people today”, “no real understanding“
TEXT: “Men are typically more suited for high-stress jobs than women”. BIAS ASPECT: Gender Bias NOTABLE WORDS/PHRASES: “Men”, “more suited”, “than women“
TEXT: “Despite their qualifications, immigrants from Asia often find it hard to get jobs in this sector”. BIAS ASPECT: Ethnicity Bias NOTABLE WORDS/PHRASES: “immigrants from Asia”, “hard to get jobs“
TEXT: “Atheists tend to have a more cynical view of life compared to those who are religious”. BIAS ASPECT: Religious Bias NOTABLE WORDS/PHRASES: “Atheists”, “more cynical view”, “compared to those who are religious“
TEXT: “People from low-income neighborhoods are more likely to engage in criminal activities”. BIAS ASPECT: Socioeconomic Bias NOTABLE WORDS/PHRASES: “low-income neighborhoods”, “more likely to engage in criminal activities“
TEXT: “Western music is inherently more sophisticated than traditional music from Africa”. BIAS ASPECT: Cultural Bias NOTABLE WORDS/PHRASES: “Western music”, “more sophisticated”, “than traditional music from Africa“
Appendix B Dataset Reviewers
B.1 Guidelines
@@
B.2 Review Guidelines Checklist
@@
B.3 Aspect Categorization